# revtex4-1Repair the float Enhancement of spin Hall angle by an order of magnitude via Cu intercalation in MoS2/CoFeB heterostructures

### v2

@@ I Introduction
Achieving high-density spin currents in materials with high spin-orbit coupling (SOC) is critical for advancing spintronic applications Hirohata et al. (2020); Soumyanarayanan et al. (2016); Ryu et al. (2020). Among the various mechanisms explored, spin pumping has emerged as a highly effective approach Tserkovnyak et al. (2002). This technique generates substantial spin current densities in high-SOC materials adjacent to a ferromagnet (FM) or ferrimagnet (FiM) with precessing magnetization. Spin pumping is particularly advantageous in heterostructures comprising metallic, insulating, or semiconducting high-SOC materials, addressing impedance mismatch and enhancing device feasibility. Traditionally, the spin Hall effect (SHE) and its Onsager reciprocal, the inverse spin Hall effect (ISHE), have been the primary mechanisms for bidirectional spin-charge conversion, driven by electron scattering in three-dimensional high-SOC materials such as heavy metals (e.g., Pt, Pd, W, and Ta), antiferromagnetic materials (e.g., IrMn, Mn2Au, Mn3Ga, Mn3Sn), and topological insulators (e.g., Bi2Se3, Bi2Te3) Hirsch (1999); Saitoh et al. (2006); Roy et al. (2021); Singh et al. (2020); Singh and Bedanta (2020); Kimata et al. (2019); Singh et al. (2019); Jamali et al. (2015). Recently, transition metal dichalcogenides (TMDs) have been identified as promising quantum materials for enhanced spin-to-charge conversion efficiency due to their high SOC Sierra et al. (2021). Among TMDs, molybdenum disulfide (MoS2) stands out for its stability, robustness, and unique electronic properties, which vary with thickness and mechanical strain Liu et al. (2020); Kumar and Panda (2022); Choi et al. (2017); Akinwande et al. (2014). In MoS2/FM heterostructures, interface effects significantly enhance global anisotropy, and MoS2 has demonstrated remarkable spin-to-charge conversion voltages and efficiencies Xie et al. (2019); Thiruvengadam et al. (2022); Mishra et al. (2024); Mendes et al. (2018); Husain et al. (2018); Bansal et al. (2019). Moreover, the efficiency of spin-to-charge conversion in FM/high-SOC systems can be further enhanced by interface engineering using intercalators such as Ag, Cu, and Ti Du et al. (2014); Longo et al. (2021); Zhang et al. (2015). These interlayers act as barriers, preserving the SOC of high-SOC materials by mitigating proximity-induced magnetic effects and tuning spin-dependent interfacial resistivity and spin injection. In this context, Cu emerges as a promising interlayer material due to its long spin diffusion length ( 350 nm), low SOC, and ability to control spin transmissivity in spintronic devices Yakata et al. (2006).
@@ II Experimental details
Thin films of MoS2(5 nm)/Cu ( nm)/CoFeB (9 nm)/Al2O3(3 nm) with varying Cu thicknesses were fabricated using a high-vacuum multi-deposition chamber (Mantis Deposition Ltd., UK) with a base pressure below 410-8 mbar. MoS2, Co40Fe40B20 (CoFeB) and Al2O3 films were deposited from commercially available stoichiometric targets. Cu and CoFeB were deposited via dc magnetron sputtering whereas MoS2 and Al2O3 were deposited by rf sputtering. The samples were grown on Si (100) substrates with a 300 nm thick SiO2 layer. Samples, labeled M1 to M7, correspond to Cu thicknesses () of 0, 0.65, 2, 3, 5, 7, and 10 nm, respectively. The reference sample, CoFeB (9 nm)/Al2O3 (3 nm), was designated as M0. Magnetization dynamics and inverse spin Hall effect (ISHE) measurements were conducted using ferromagnetic resonance (FMR). Saturation magnetization was measured with a SQUID-based magnetometer (MPMS 3, Quantum Design). Room-temperature longitudinal magneto-optic Kerr effect (MOKE) microscopy and magnetometry were employed to study magnetic domains and hysteresis loops, respectively (see Fig. S7, SI). Further details of the experimental protocols can be found in the supplementary information (SI).
@@ III Results and discussion
The presence of MoS2 in the samples was confirmed through laser Raman spectroscopy, while the thickness and surface roughness were estimated using X-ray reflectivity (XRR). Detailed structural information obtained from Raman spectroscopy and XRR is provided in the Supplementary Information (SI). The Gilbert damping parameter () was extracted by fitting the resonance field () and linewidth () data obtained from ferromagnetic resonance (FMR) spectroscopy, as described in the SI. The values (10-3) for samples M1–M7 are listed in Table I and are higher than that of the reference sample M0 (8.20.1), indicating potential spin pumping effects. However, other contributing mechanisms cannot be ruled out. To verify spin pumping, ISHE measurements were performed. Figure 1(a) shows the variation in measured ISHE voltage () as a function of applied field (H) for sample M5 at in-plane angles () of 0∘ and 180∘. The symmetric (Vsym) and antisymmetric (Vasym) components of were separated by fitting the data to the following Lorentzian equation, Conca et al. (2017),
@@
The initial decrease in can be attributed to increased roughness in the ultrathin Cu layer (0.65 nm). At such small thicknesses, finite size effects become significant, primarily due to enhanced electron scattering at surfaces and grain boundaries, which raises the overall resistivity Zhou and Gall (2018). The resistance of the samples varies from 3620.45 to 260.09 from sample M1 to M7 which indicates the shunting effect due to Cu. This indicates that at low Cu thickness, electrons experience frequent scattering events, including spin-flip scattering, which disrupts spin alignment and reduces spin transport efficiency, thereby decreasing the spin Hall angle. With increasing Cu thickness, the Cu layer starts to act as a good spin transmitter with minimal dissipation and reduced spin flipping, preserving the spin polarization. This leads to an increase in the spin Hall angle. At 5 nm, the Cu layer becomes sufficiently thick to support spin accumulation. Spin accumulation refers to the buildup of spin-polarized electrons at the interfaces, enabled by a thickness that allows spins to diffuse without substantial spin-flip scattering. This suggests that the spin Hall angle of the Cu interlayer generates a positive voltage response, reducing the net spin Hall angle Wang et al. (2014). In the MoS2/CoFeB bilayer structure, the direct interface likely causes significant spin dissipation and the formation of mixed electronic states, both of which impede efficient spin accumulation. By introducing a Cu interlayer, these effects are mitigated, creating a cleaner pathway for spin currents to reach the MoS2 layer and improving overall spin transport efficiency.
To further validate the decoupling in the MoS2/Cu/CoFeB trilayer, we conducted MOKE microscopy. The domain images in Fig. 3 (a–e) reveal the isotropic magnetic behavior of the reference CoFeB single-layer sample (M0). These findings suggest that the reference CoFeB layer exhibits magnetic isotropy within the film plane, a characteristic typical of amorphous ferromagnets, where the absence of long-range structural order minimizes magnetocrystalline anisotropy Xie et al. (2020); Nayar and Batra (1990). In contrast, domain images of M1 (MoS2/CoFeB) at 0∘ and 180∘, shown in Fig. 3(f) and Fig. 3(j), display larger domains, indicative of magnetization reversal driven by domain wall nucleation and motion Chowdhury and Bedanta (2014). These observations suggest that MoS2 introduces significant anisotropy in the CoFeB layer, consistent with earlier findings Thiruvengadam et al. (2022). However, when a Cu spacer layer is introduced between MoS2 and CoFeB, the domain images for M5 (Fig. 3(k–o)) closely resemble those of the reference CoFeB layer, indicating that the CoFeB magnetic microstructure is no longer influenced by MoS2. This confirms that the Cu interlayer effectively mitigates proximity-induced effects at the MoS2/CoFeB interface. By decoupling the two layers, the Cu spacer facilitates improved spin transport and conversion efficiency.
@@ IV Theoretical calculations
To understand the role of the Cu spacer layer in the spin-pumping we performed first-principles based calculation to estimate the spin Hall angles (SHA) in MoS2/CoFeB and MoS2/Cu/CoFeB systems within the framework of Density Functional Theory (DFT). The detailed methodology of these calculations is mentioned in the SI. The spin Hall angle is given by Qu et al. (2014),
@@ V Conclusion
We achieved a substantial enhancement in spin-to-charge conversion efficiency, with an increase of an order of magnitude, through Cu intercalation in the MoS2/CoFeB system. Domain imaging confirmed that Cu effectively decouples MoS2 from the proximity effects of CoFeB. The experimental findings are corroborated by theoretical calculations using the DFT-WANNIER90 framework, which reveal significant improvements in spin Berry curvature, spin Hall conductivity, and magnetic anisotropy energy. In summary, Cu intercalation preserves the intrinsic SOC of MoS2 by mitigating magnetic exchange interactions with CoFeB, resulting in a marked enhancement of the spin Hall effect. These results highlight the potential of intercalation-based interface engineering strategies to enable the development of highly efficient spintronic devices.

### v3

# Enhancement of spin Hall angle by an order of magnitude via Cu intercalation in MoS2/CoFeB heterostructures

@@ I Introduction
Achieving high-density spin currents in materials with high spin-orbit coupling (SOC) is critical for advancing spintronic applications Hirohata et al. (2020); Soumyanarayanan et al. (2016); Ryu et al. (2020). Among the various mechanisms explored, spin pumping has emerged as a highly effective approach Tserkovnyak et al. (2002). This technique generates substantial spin current densities in high-SOC materials adjacent to a ferromagnet (FM) or ferrimagnet (FiM) with precessing magnetization. Spin pumping is particularly advantageous in heterostructures comprising metallic, insulating, or semiconducting high-SOC materials, addressing impedance mismatch and enhancing device feasibility. Traditionally, the spin Hall effect (SHE) and its Onsager reciprocal, the inverse spin Hall effect (ISHE), have been the primary mechanisms for bidirectional spin-charge conversion, driven by electron scattering in three-dimensional high-SOC materials such as heavy metals (e.g., Pt, Pd, W, and Ta), antiferromagnetic materials (e.g., IrMn, Mn2Au, Mn3Ga, Mn3Sn), and topological insulators (e.g., Bi2Se3, Bi2Te3) Hirsch (1999); Saitoh et al. (2006); Roy et al. (2021); Singh et al. (2020); Singh and Bedanta (2020); Kimata et al. (2019); Singh et al. (2019); Jamali et al. (2015). Recently, transition metal dichalcogenides (TMDs) have been identified as promising quantum materials for enhanced spin-to-charge conversion efficiency due to their high SOC Sierra et al. (2021). Among TMDs, molybdenum disulfide (MoS2) stands out for its stability, robustness, and unique electronic properties, which vary with thickness and mechanical strain Liu et al. (2020); Kumar and Panda (2022); Choi et al. (2017); Akinwande et al. (2014). In MoS2/FM heterostructures, interface effects significantly enhance global anisotropy, and MoS2 has demonstrated remarkable spin-to-charge conversion voltages and efficiencies Xie et al. (2019); Thiruvengadam et al. (2022); Mishra et al. (2024); Mendes et al. (2018); Husain et al. (2018); Bansal et al. (2019). Furthermore, several mechanisms have been suggested to improve the efficiency of spin-to-charge conversion in FM/high-SOC systems Du et al. (2014); Longo et al. (2021); Obstbaum et al. (2016); Chen et al. (2024); Choi et al. (2024). Among them, the utilization of interlayers has been an effective approach. These interlayers act as barriers, preserving the SOC of high-SOC materials by mitigating proximity-induced magnetic effects and tuning spin-dependent interfacial resistivity and spin injection. In this context, Cu emerges as a promising interlayer material due to its long spin diffusion length ( 350 nm), low SOC, and ability to control spin transmissivity in spintronic devices Yakata et al. (2006).
@@ II Experimental details
Thin films of MoS2(5 nm)/Cu ( nm)/CoFeB (9 nm)/Al2O3(3 nm) with varying Cu thicknesses were fabricated using a high-vacuum multi-deposition chamber (Mantis Deposition Ltd., UK) with a base pressure below 4 10-8 mbar. MoS2, Co40Fe40B20 (CoFeB) and Al2O3 films were deposited from commercially available stoichiometric targets. Cu and CoFeB were deposited via dc magnetron sputtering whereas MoS2 and Al2O3 were deposited by rf sputtering. The samples were grown on Si (100) substrates with a 300 nm thick SiO2 layer. Samples, labeled M1 to M7, correspond to Cu thicknesses () of 0, 0.65, 2, 3, 5, 7, and 10 nm, respectively. The reference sample, CoFeB (9 nm)/Al2O3 (3 nm), was designated as M0. Magnetization dynamics and inverse spin Hall effect (ISHE) measurements were conducted using ferromagnetic resonance (FMR). Saturation magnetization was measured with a SQUID-based magnetometer (MPMS 3, Quantum Design). Room-temperature longitudinal magneto-optic Kerr effect (MOKE) microscopy and magnetometry were employed to study magnetic domains and hysteresis loops, respectively (see Fig. S7 of Supplemental Material SM_ )). Further details of the experimental protocols can be found in the Supplemental Material SM_ .
@@ III Results and discussion
The presence of MoS2 in the samples was confirmed through laser Raman spectroscopy, while the thickness and surface roughness were estimated using X-ray reflectivity (XRR). Detailed structural information obtained from Raman spectroscopy and XRR is provided in the Supplemental Material SM_ . The Gilbert damping parameter () was extracted by fitting the resonance field () and linewidth () data obtained from ferromagnetic resonance (FMR) spectroscopy, as described in the Supplemental Material SM_ . The values (10-3) for samples M1–M7 are listed in Table I and are higher than that of the reference sample M0 (8.20.1), indicating potential spin pumping effects. However, other contributing mechanisms cannot be ruled out. To verify spin pumping, ISHE measurements were performed. In-plane angle-dependent ISHE measurements are particularly advantageous because they provide a direct, controlled, and quantitative probe of spin-charge conversion with high sensitivity. Figure 1(a) shows the variation in measured ISHE voltage () as a function of applied field (H) for sample M5 at in-plane angles () of 0∘ and 180∘. The symmetric (Vsym) and antisymmetric (Vasym) components of were separated by fitting the data to the following Lorentzian equation, Conca et al. (2017),
@@
The resistance of the heterostructures decreases gradually due to Cu intercalation and varies from 3620.45 to 260.09 from sample M1 to M7. With increasing Cu thickness, the Cu layer starts to act as a good spin transmitter with minimal dissipation and reduced spin flipping, preserving the spin polarization. This leads to an increase in the spin Hall angle. At 5 nm, the Cu layer becomes sufficiently thick to support spin accumulation. Spin accumulation refers to the buildup of spin-polarized electrons at the interfaces, enabled by a thickness that allows spins to diffuse without substantial spin-flip scattering. This suggests that the spin Hall angle of the Cu interlayer generates a positive voltage response, reducing the net spin Hall angle Wang et al. (2014). In the MoS2/CoFeB bilayer structure, the direct interface likely causes significant spin dissipation and the formation of mixed electronic states, both of which impede efficient spin accumulation. By introducing a Cu interlayer, these effects are mitigated, creating a cleaner pathway for spin currents to reach the MoS2 layer and improving overall spin transport efficiency.
To further validate the interfacial decoupling in the MoS2/Cu/CoFeB trilayer, we performed MOKE microscopy. When MoS2 is directly interfaced with CoFeB, we observe larger magnetic domains, suggesting that magnetization reversal is dominated by domain wall nucleation and motion Chowdhury and Bedanta (2014). This is consistent with interfacial hybridization effects reported in earlier findings Thiruvengadam et al. (2022). This can be observed from the domain image of M1 (MoS2/CoFeB) at 0∘, shown in Fig. 3(b). However, when Cu is intercalated between MoS2 and CoFeB, it displays fuzzy domains shown in Fig. 3 (c). This is similar to that in the reference bare CoFeB layer (Fig. 3 (a)).
It is to be noted that all our samples, with or without a Cu insertion layer, exhibit in-plane magnetic anisotropy. The effective anisotropy energy density has been quantified via SQUID magnetometry (see Supplemental Material SM_ ). The for samples M0, M1 and M5 are (6.970.28)106, (5.670.21)106 and (6.650.32)106erg/cc, respectively. Therefore, the effective anisotropy is found to be enhanced with Cu intercalation. Similar trend has also been found from the DFT calculations (to be discussed later).
@@ IV Theoretical calculations
To understand the role of the Cu spacer layer in the spin-pumping we performed first-principles based calculation to estimate the spin Hall angles (SHA) in MoS2/CoFeB and MoS2/Cu/CoFeB systems within the framework of Density Functional Theory (DFT). The detailed methodology of these calculations is mentioned in the Supplemental Material SM_ . The spin Hall angle is given by Qu et al. (2014),
@@ V Conclusion
We achieved a substantial enhancement in spin-to-charge conversion efficiency, with an increase of an order of magnitude, through Cu intercalation in the MoS2/CoFeB system. Domain imaging confirmed that Cu effectively decouples MoS2 from the proximity effects of CoFeB. The experimental findings are corroborated by theoretical calculations using the DFT-WANNIER90 framework, which reveal significant improvements in spin Berry curvature, spin Hall conductivity, and magnetic anisotropy energy. In summary, Cu intercalation preserves the intrinsic SOC of MoS2 by mitigating magnetic exchange interactions with CoFeB, resulting in a marked enhancement of the spin Hall effect. This approach can further be explored to enhance the spin-to-charge conversion efficiency in other TMD materials like WSe2, MoSe2, MoTe2 etc. These results highlight the potential of intercalation-based interface engineering strategies to enable the development of highly efficient spintronic devices.